OpenStudy (ashking1):

What are the solutions for x in the proportion (5x-6)/(x)=(2x)/(2)

OpenStudy (mertsj):

Did you cross multiply?

OpenStudy (mertsj):

\[\frac{5x-6}{x}=\frac{2x}{2}\]

OpenStudy (pfenn1):

\[\frac{5x-6}{x}=\frac{2x}{2}\]You can simplify \[\frac{2x}{2}=x\]Then multiply both sides of the equation by x\[5x-6=x^2\]

OpenStudy (ashking1):

x = 2 and x = 3 x = -12 and x = -8 x = 12 and x = 8 x = -2 and x = -3 these are the choices

OpenStudy (pfenn1):

Factor or use the quadratic formula to solve.

OpenStudy (ashking1):

ok i got it thanks

OpenStudy (pfenn1):

You're welcome.